OpenStudy (anonymous):

simplify 12a^3b^2 divided by 4ab^-2 show all working out

OpenStudy (anonymous):

\[(12a^{3}b ^{2})\div(4ab ^{-2}) is coeerct?\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

12:4 = 3\[a^{2}\div a ^{1}=a ^{2-1} =a ^{1} =a\]

OpenStudy (anonymous):

\[b^{2} \div b ^{-2}=b ^{2-(-2)} =b ^{4}\]

OpenStudy (anonymous):

\[3ab ^{4}\]

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

a^3 not a^2

OpenStudy (anonymous):

give me more time I want double check

OpenStudy (anonymous):

take all the time you need i am greatful for your help

OpenStudy (anonymous):

\[a^{3} \div a=a ^{3-1}=a^{2}\]

OpenStudy (anonymous):

\[3a ^{2}b ^{4}\] here correct answer,I double check

OpenStudy (anonymous):

any question?

OpenStudy (anonymous):

take time look the answer ,I have to go is late

OpenStudy (anonymous):

given that E=mc^2, find c if m=0.05 and E=4.5*10^15 (A) 1.5*10^7 (B) 3*10^8 (C) 1.0*10^15 (D) 2*10^16 or none of the above